Third Bank has the following balance sheet (in millions of dollars) with the risk weights in parentheses. Assets Liabilities and equ Cash (096 Interbank deposits with AA rated banks (20%) Standard residential mortgages non- insured with LVR of 85% (50% Business loans to BB rated borrowers (100%) Total assets 20 Deposits 175 25 Subordinated debt (5 years) 70 Cumulative preference shares 70 Common equity (Tier 1) 185 Total liabilities and equit Tier 2 capita er 1 185 In addition, the bank has $30 million in performance-related standby letters of credit (SLCs) $40 million in two-year forward FX contracts that are currently in the money by $1 million, and $300 million in six-year interest rate swaps that are currently out of the money by $2 million (a) What are the risk-adjusted on-balance-sheet assets of the bank as defined under the Basel Accord? Cash Interbank deposits Mortgage loans Business loans Total risk-adjusted assets Calculation 0 20 0.2 25 0.5*70 1*70 185 Risk Weighted value 0 5 35 70 (b) Discuss briefly under which circumstances the two-year forward FX contracts and the interest rate swaps introduce an additional risk for the bank. (c) Presume that the off-balance sheet risk is evaluated at $19.25 million. What is the total amount of risk-adjusted on- and off-balance-sheet assets?
Answer #1

1) Risk-adjusted assets:

Cash

0 * 20

0

Interbank deposits

0.20 * 25

5

Mortgage loans

0.50 * 70

35

Business loans

1.00 * 70

70

Total risk-adjusted assets

110

110

2) It will depend on whether the bank used these derivatives as a hedge for insuring against existing on-balance sheet risk or as speculative investment. In the first scenario the risk will be reduced for the bank, and in the second scenario there will be an extra risk. For eample: When the bank had FX exposure on the balance sheet and an exactly opposite FX exposure with the off- balance derivative, it will not be exposed to FX risk anymore. But when the FX forwards are not a counter position to existing assets (or liabilities), the movements of FX can push them out off money, thus causing huge losses for the bank

3) No, because bank lacks the sufficient capital to meet the regulatory requirements of capital.

Common equity Tier 1 CAR = 1.54% (=2/129.25)

The interest is less than 7% as required for Basel III and APRA

All Tier 1 capital is only $7 million, producing a Tier 1 CAR = 5.42% (=7/129.25)

The interest is less than 8.5% as required for Basel III and APRA

Total CAR = 7.73% (=10/129.25)

The interest is less 10.5% as required for Basel III and APRA

Thus, the bank fails all three CARs.

If it increases the sufficient equity to meet the common equity Tier 1 CAR, in that case it will also satisfy all i.e. CARs: 5.42+5.46 > 10.5 > 8.5